# Universal Jailbreak Suffixes Are Strong Attention Hijackers

### v1

@@ Abstract
We study suffix-based jailbreaks—a powerful family of attacks against large language models (LLMs) that optimize adversarial suffixes to circumvent safety alignment. Focusing on the widely used foundational GCG attack Zou et al. (2023b), we observe that suffixes vary in efficacy: some markedly more universal—generalizing to many unseen harmful instructions—than others. We first show that GCG’s effectiveness is driven by a shallow, critical mechanism, built on the information flow from the adversarial suffix to the final chat template tokens before generation. Quantifying the dominance of this mechanism during generation, we find GCG irregularly and aggressively hijacks the contextualization process. Crucially, we tie hijacking to the universality phenomenon, with more universal suffixes being stronger hijackers. Subsequently, we show that these insights have practical implications: GCG universality can be efficiently enhanced (up to 5 in some cases) at no additional computational cost, and can also be surgically mitigated, at least halving attack success with minimal utility loss. We release our code and data at github.com/matanbt/interp-jailbreak.
@@ 1 Introduction
The rapid adoption of Transformer-based large language models (LLMs) has raised concerns about misuse, including harmful content generation. While safety alignment—fine-tuning LLMs to prevent such outputs—has become standard Bai et al. (2022); Rafailov et al. (2024), these safeguards remain vulnerable to jailbreak attacks that bypass alignment by manipulating prompts Zou et al. (2023b); Wei et al. (2023); Chao et al. (2024b). Suffix-based jailbreaks such as GCG (Zou et al., 2023b) append a short, unintelligible sequence to a harmful instruction, reliably causing model compliance. This family of attacks, popularized and underpinned by GCG, is now a standard tool for automated red-teaming Chao et al. (2024a) and represents a powerful class of jailbreaks Sadasivan et al. (2024); Thompson and Sklar (2024); Hayase et al. (2024); Andriushchenko et al. (2025). Remarkably, many such suffixes exhibit universality, generalizing to diverse unseen instructions Zou et al. (2023b), even, as we show, when optimized for a single harmful behavior (§2).
@@
Recent work turned to interpretability to understand safeguard and jailbreak mechanisms Zou et al. (2023a); Ball et al. (2024); Kirch et al. (2024); Arditi et al. (2024); Jain et al. (2024); Lee et al. (2024); Li et al. (2025). These analyses mostly focus on the internal representation of the last token position before generation, compare the representation for harmful and benign prompts to extract harmfulness-related directions, show jailbreaks shift away from these directions, and use them to steer the model behavior through low-rank modifications of its representations.
@@
Yet, the effective internal mechanisms employed by jailbreaks—and, in particular, by suffix-based jailbreaks— remain far from being fully understood. Notably, the common focus of the last token position in jailbreak analyses lacks systematic localization; nor is it clear what characterizes the mechanism enabling the jailbreak and preceding the (dis)appearance of previously found directions, what differentiates more universal suffixes, or whether such insights can be used in practice to improve jailbreak efficacy or mitigation.
@@
In this paper, we systematically localize (§4), surgically investigate (§5–6), and practically utilize (§7) the underlying mechanism of suffix-based jailbreaks, focusing on the common, representative GCG attack Zou et al. (2023b).
@@
First, we localize the jailbreak mechanism to the critical information flow originating from the adversarial suffix ( ; Fig. 1), finding it to be shallow—concentrated in the chat template tokens immediately preceding generation ( ; Fig. 1). We establish that this information flow is both invariably necessary and mostly sufficient for the jailbreak to succeed. These findings justify the prior focus of jailbreak interpretability on the last token position ( ), and align with recent research Qi et al. (2025), which shows that current safety alignment methods shallowly focus on the first few generated tokens.
@@
Second, we examine this shallow mechanism, identifying a phenomenon common in GCG suffixes, and rare among other suffix distributions, which we term hijacking. Building on Kobayashi et al. (2021); Mickus et al. (2022), we quantify ’s dominance in ’s contextualization. We find that GCG suffixes consistently attain exceptionally high dominance, effectively hijacking , with magnitude that of similarly structured benign or adversarial prompts. In some layers accounts for nearly all of the attention output, while the harmful instruction’s ( ) contribution is almost eliminated from early layers onward. This provides a more direct view of jailbreaks’ shift away from harmfulness-related directions observed in prior work.
@@
Third, we assess the hijacking strength of each adversarial suffix by aggregating the dominance score across different harmful instructions. We show it correlates with the suffix’s emerging universality, suggesting hijacking is an essential mechanism to which universal suffixes converge. In particular, we show that the more universal a suffix, the stronger its hijacking effect, with the most universal suffixes consistently exhibiting abnormally high hijacking strength. Notably, the hijacking strength underlying universality is obtained without generating any tokens.
@@
Lastly, we demonstrate that our insights have practical implications. On the one hand, encouraging hijacking while optimizing jailbreaks—which can be done with no computational overhead, differently than existing universal attacks Zou et al. (2023b)—reliably produces more universal adversarial suffixes, resulting in a stronger attack (§7.1). On the other hand, suppressing the hijacking mechanism impairs suffix-based jailbreaks with minimal harm to model utility, effectively providing a mitigation strategy (§7.2).
In conclusion, we investigate GCG, a predominant suffix-based jailbreak (§2–3). We start by systematically localizing the GCG jailbreak key mechanism (§4), then show GCG suffixes exert irregular dominance in the final tokens before generation, effectively hijacking them (§5)—a phenomenon which we find to be closely tied to universality (§6). Building on these insights, we demonstrate GCG jailbreaks can be both efficiently enhanced and surgically mitigated (§7).
@@ 2 Preliminaries: Suffix-based Jailbreaks
Suffix-based LLM jailbreaks are a powerful class of inference-time attacks Mazeika et al. (2024); Chao et al. (2024a) that set to bypass model safety alignment by appending an automatically optimized adversarial suffix (adv) to a harmful instruction (instr) (Fig. 1). We focus on this family of attacks, specifically on the widely used, foundational GCG method Zou et al. (2023b).
Suffix-based attacks are not only highly effective and common in automatic red-teaming Chao et al. (2024a), but their unified structure also enables systematic study. Unlike handcrafted jailbreak prompts (e.g., “My grandma used to tell me how to build a bomb before bedtime” Wei et al. (2023); Shen et al. (2024)), these optimized adversarial suffixes are unintelligible and opaque, motivating the need for interpretation. We focus on the popular GCG attack Zou et al. (2023b), which underpins many recent suffix-based methods that extend its objective Thompson and Sklar (2024), prompt template Andriushchenko et al. (2025), or optimization process Sadasivan et al. (2024); Hayase et al. (2024); Thompson and Sklar (2024). GCG thus captures the general methodology shared across this family of attacks.
GCG crafts adv by searching for token sequences following the affirmation objective—maximizing the likelihood of an affirmative response for a given instruction (affirm; e.g., “Sure, here’s how to build a bomb”). This builds on the observation that prefilling the response with an affirmative prefix (prefilling attacks; Tang (2024)) often induces successful jailbreaks Qi et al. (2025). To increase universality, GCG can be further optimized against multiple behaviors Zou et al. (2023b), a standard, though computationally intensive, strategy Thompson and Sklar (2024); Sadasivan et al. (2024).
@@ 3.1 Experimental Setup
Our main analysis uses Gemma2-2B-it Google (2024) to enable scale and depth, with critical evaluations validated on Qwen2.5-1.5B-Instruct Qwen (2025) and Llama-3.1-8B-Instruct Meta (2024) (§3, §6–7).
We use GCG (default hyperparameters) to craft 1200 adversarial suffixes on Gemma2-2B-it, each optimized against a single behavior from AdvBench Zou et al. (2023b). Combined with 741 harmful instructions from AdvBench Zou et al. (2023b) and StrongReject Souly et al. (2024), these yield nearly 900K GCG jailbreak prompts of varying success, used throughout the paper. We also generate 100 GCG suffixes each for Qwen2.5 and Llama3.1 for additional evaluation. We defer more technical details to the appendix (App. A.1).
We measure jailbreak success using StrongReject’s fine-tuned classifier Souly et al. (2024), which assigns a grade in , with higher values indicating more effective jailbreaks. We label attack samples as successful (), failed (), or borderline (otherwise), based on the classifier’s grading of harmful response quality. A suffix’s universality score is defined as its success rate across the full set of harmful instructions.
@@ 3.2 Characterizing GCG Suffixes
We analyze over 1K single-instruction GCG suffixes on Gemma2 (for Qwen2.5 and Llama3.1 see App. B.1) and reveal that (a) GCG suffixes show varying level of efficacy, and (b) they often generalize beyond their explicit affirmation objective.
First, single-instruction GCG suffixes often generalize beyond their target instruction, exhibiting varying universality, a phenomenon also noted in contemporary work by Huang et al. (2025). As Fig. 2(a) shows, most suffixes jailbreak over 2% of the tested harmful instructions, with the strongest succeeding in 20–60% of cases.
@@
These observations motivate our central question: what underlying mechanisms enable the effectiveness of different GCG’s suffixes, and particularly the emergent strong, universal suffixes?
@@ 4.1 Localizing the Critical Information Flow
Aiming to localize the critical information flow from the adversarial suffix (adv), we perform attention knockout Geva et al. (2023), as it is the sole component enabling information to transfer across token representations Elhage et al. (2022).
Experimental setting. We sample 1K successful jailbreaks across suffixes of diverse universality, and perform attention knockout on each edge departing from adv to following token subsequences (Fig. 1), by masking the edge’s attention (i.e., setting its attention logits to , in all layers). Then, we measure the Jailbreak Flip Rate (JFR): the fraction of attacks that knockout flips from success to failure. Higher JFR indicates greater edge importance in generation supporting the jailbreak.
Knockout adv. Fig. 3(a) shows the JFR of different suffixes, for each edge. We find advchat to be overwhelmingly critical for the jailbreak; knocking out advchat consistently fails the attack (causes refusal), whereas other edges (e.g., advaffirm) only occasionally do so. Notably, the removal of advchat, prevents the jailbreak from manipulating the model into starting the generation with an affirmative token (e.g., “Sure”), which, as observed by Qi et al. (2025), may single-handedly fail the jailbreak. To rule out this case, we perform a series of ablation studies on advchat’s knockout, forcing the generation to start with various dummy tokens (e.g., white spaces, additional sequence of chat tokens, and random punctuation), as well as an affirmative token (e.g., Sure), and find that the strong presented trend is kept (JFR ; Fig. 10, App. B.2).
Knockout adv, under prefilling. Given the former results (Fig. 3(a)), it is possible that the role of advchat is primarily to supply the affirmative response prefix, thus, failing the affirmation implies failing the jailbreak. In what follows, we rule this out. We repeat the knockout, this time applying it under prefilling, that is, starting the generation after an affirmative response prefix (§3.1). As Fig. 3(b) demonstrates, advchat still has the highest JFR among edges, with most suffixes having JFR (i.e., generally, this edge’s knockout mainly fails prefilled attack). This shows the suffixes’ critical role extends beyond naïvely inducing affirmation.
@@ 4.2 Restoring Jailbreak via Shallow Patching
Experiment setting. To isolate the causal effect of the information flow to chat on the jailbreak behavior, we perform a causal mediated analysis Vig et al. (2020), by patching chat’s attention output activations Wang et al. (2023); Zhang and Nanda (2024).111Borrowing Vig et al. (2020)’s terms, we test the indirect effect of the input prompt (specifically, of adv) on the jailbreak behavior (i.e., a property of the output), while viewing chat as the inspected mediator. For 300 instruction-matched pairs of failed attacks (either a random or GCG suffixes) and successful attacks, we take the failed sample and patch its attention outputs at chat (all layers) with those from the successful counterpart. Namely, we form each patched example by retaining the failed prompt (including adv) and injecting the successful sample’s chat activations; if this restores the jailbreak, we attribute it to the transferred advchat pathway. To control the patch depth, we incrementally extend the patch to i tokens after chat (denoted chat+i).
Patching chat+i. Fig. 4 shows that patching only chat restores the jailbreak behavior in the majority of cases. This effect is slightly enhanced when increasing the depth of the patched token subsequence into the generation (i.e., i in chat+i), with the majority of the advancement taking effect only a few tokens deep, suggesting the attacks’ key mechanism is shallow.

### v2

@@ Abstract
We study suffix-based jailbreaks—a powerful family of attacks against large language models (LLMs) that optimize adversarial suffixes to circumvent safety alignment. Focusing on the widely used foundational GCG attack zou2023gcg-universaltransferableadversarialattacks, we observe that suffixes vary in efficacy: some are markedly more universal—generalizing to many unseen harmful instructions—than others. We first show that a shallow, critical mechanism drives GCG’s effectiveness. This mechanism builds on the information flow from the adversarial suffix to the final chat template tokens before generation. Quantifying the dominance of this mechanism during generation, we find GCG irregularly and aggressively hijacks the contextualization process. Crucially, we tie hijacking to the universality phenomenon, with more universal suffixes being stronger hijackers. Subsequently, we show that these insights have practical implications: GCG’s universality can be efficiently enhanced (up to 5 in some cases) at no additional computational cost, and can also be surgically mitigated, at least halving the attack’s success with minimal utility loss.111We release our code and data at https://github.com/matanbt/interp-jailbreak.
@@ 1 Introduction
The rapid adoption of Transformer-based large language models (LLMs) has raised concerns about misuse, including harmful content generation. While safety alignment—fine-tuning LLMs to prevent such outputs—has become a common practice bai2022traininghelpfulharmlessassistant; rafailov2024directpreferenceoptimizationlanguage, these safeguards remain vulnerable to jailbreak attacks that bypass alignment by manipulating prompts zou2023gcg-universaltransferableadversarialattacks; wei2023jailbrokendoesllmsafety; chao2024PAIR-jailbreakingblackboxlarge. Suffix-based jailbreaks such as GCG (zou2023gcg-universaltransferableadversarialattacks) append a short, unintelligible sequence to a harmful instruction, reliably causing model compliance. This family of attacks, popularized and underpinned by GCG, is now a ubiquitous tool for automated red-teaming chao2024jailbreakbenchopenrobustnessbenchmark and represents a powerful class of jailbreaks sadasivan2024fastadversarialattackslanguage; thompson2024flrtfluentstudentteacherredteaming; hayase2024querybasedadversarialpromptgeneration; andriushchenko2025PRSjailbreakingleadingsafetyalignedllms. Remarkably, many such suffixes exhibit universality on the targeted model, generalizing to diverse unseen instructions zou2023gcg-universaltransferableadversarialattacks, even, as we show, when optimized for a single harmful behavior (§3).
@@
Recent work turned to interpretability to analyze safeguard and jailbreak mechanisms zou2023representationengineeringtopdownapproach; ballUnderstandingJailbreakSuccess2024; kirchWhatFeaturesPrompts2024; arditiRefusalLanguageModels2024; jainWhatMakesBreaks2024a; leeMechanisticUnderstandingAlignment2024; liRevisitingJailbreakingLarge2024; leong2025safeguardedshipsrunaground. These analyses mostly focus on the internal representation of the last token position before generation, compare harmful and benign prompts’ representations to extract harmfulness-related directions, show jailbreaks shift away from these directions, and use them to steer the model behavior through modifications of its representations. Other work suggested existing safety alignment methods are shallow—i.e., their effect concentrates on the first few generated tokens qi2024safetyalignmentjusttokens.
@@
Yet, the effective internal mechanisms employed by jailbreaks—and, in particular, by suffix-based jailbreaks—remain far from being fully understood. Notably, the common focus on the last token position in jailbreak analyses lacks systematic localization; it is unclear whether, and if so how, jailbreaks internally exploit the shallowness of safety alignment; nor is it clear what characterizes the mechanism enabling the jailbreak and preceding the (dis)appearance of previously found directions, what differentiates more universal suffixes, or whether such insights can be used in practice to improve jailbreak efficacy or mitigation.
@@
In this paper, we address these gaps by investigating suffix-based jailbreaks (§2–3); we systematically localize (§4), surgically investigate (§5–6), and practically utilize (§7) suffix-based jailbreaks’ mechanism, focusing on the common, representative GCG attack zou2023gcg-universaltransferableadversarialattacks.
@@
First, we localize the jailbreak mechanism to the critical information flow from the adversarial suffix ( ; Fig. 1), finding it to be shallow—concentrated in the internal representation of the chat template tokens preceding generation ( ; Fig. 1). We establish that this information flow is both invariably necessary and mostly sufficient for the jailbreak to succeed. These findings justify the prior focus of jailbreak interpretability on the last token position ( ; e.g., kirchWhatFeaturesPrompts2024; ballUnderstandingJailbreakSuccess2024).
@@
Second, we examine this shallow jailbreak mechanism, identifying a phenomenon common in GCG suffixes, and rare among other suffix distributions, which we term hijacking. Building on the work of kobayashiIncorporatingResidualNormalization2021 and mickusHowDissectMuppet2022, we quantify the contribution dominance of the different input subsequences (e.g., , ) to ’s representation. We find that GCG suffixes consistently attain exceptionally high dominance, effectively hijacking ’s contextualization, to an extent that surpasses even similarly structured benign or adversarial prompts. Namely, for GCG prompts, accounts for nearly all of the attention output in some layers, while the harmful instruction’s ( ) contribution is almost eliminated from early layers onward. This analysis provides a more direct view of jailbreaks’ shift away from harmfulness-related directions observed in prior works, and shows how suffix-based jailbreaks mechanistically exploit the vulnerability of shallow safety alignment qi2024safetyalignmentjusttokens; leong2025safeguardedshipsrunaground.
@@
Third, we assess the hijacking strength of each adversarial suffix by aggregating the dominance score across different harmful instructions. We show that, for various models, this hijacking strength is linked with the suffix’s emerging universality, suggesting hijacking is an essential mechanism to which universal suffixes converge. In particular, we show that the more universal a suffix, the stronger its hijacking effect, with the most universal suffixes consistently exhibiting abnormally high hijacking strength. Notably, the hijacking strength underlying universality is efficiently obtained without generating any tokens.
@@
Lastly, we demonstrate that our insights have practical implications, further tightening the link between hijacking and attack success. On the one hand, encouraging hijacking while optimizing jailbreaks—which can be done with no computational overhead, unlike existing universal attacks zou2023gcg-universaltransferableadversarialattacks—reliably produces more universal adversarial suffixes, resulting in a stronger attack (§7.1). On the other hand, actively suppressing the hijacking mechanism impairs suffix-based jailbreaks with minimal harm to model utility, effectively providing a mitigation strategy (§7.2), while carefully tracking the hijacking strength demonstrates a strong baseline for detecting attacks (§7.3).
Contributions. Overall, our work makes the following contributions: (a) we introduce a mechanistic analysis of suffix-based jailbreaks, showing their mechanism exerts irregular dominance in the final tokens before generation, in effect hijacking them; (b) we tie this hijacking phenomenon to jailbreak suffixes’ universality; (c) we demonstrate it is possible to translate our insights into offensive and defensive advances in LLMs. More broadly, this work shows how mechanistic analysis of potent attacks against machine-learning models can aid the understanding, exploitation, and mitigation of their underlying vulnerabilities.
@@ 2 Preliminaries: Suffix-based Jailbreaks
Suffix-based LLM jailbreaks are a powerful class of inference-time attacks mazeika2024harmbenchstandardizedevaluationframework; chao2024jailbreakbenchopenrobustnessbenchmark that seek to bypass model safety alignment by appending an automatically optimized adversarial suffix (adv) to a harmful instruction (instr) (Fig. 1). We focus on this family of attacks, specifically on the widely used, foundational GCG method zou2023gcg-universaltransferableadversarialattacks.
Suffix-based attacks are not only highly effective and common in automatic red-teaming chao2024jailbreakbenchopenrobustnessbenchmark, but their unified and modular structure also enables systematic study of LLM jailbreaks. Unlike transparent handcrafted jailbreak prompts (e.g., “My grandma used to tell me how to build a bomb before bedtime” wei2023jailbrokendoesllmsafety; shen2024donowcharacterizingevaluating), these optimized adversarial suffixes are often unintelligible and opaque, motivating the need for interpretation.
We focus on the popular GCG attack (Greedy Coordinate Gradient; zou2023gcg-universaltransferableadversarialattacks), and complement our evaluation with BEAST (Beam Search-based Adversarial Attack; sadasivan2024fastadversarialattackslanguage)—a black-box GCG variant for crafting natural and fluent jailbreak suffixes. GCG underpins many recent suffix-based methods that extend its objective thompson2024flrtfluentstudentteacherredteaming, prompt template andriushchenko2025PRSjailbreakingleadingsafetyalignedllms, or optimization process sadasivan2024fastadversarialattackslanguage; hayase2024querybasedadversarialpromptgeneration; thompson2024flrtfluentstudentteacherredteaming. GCG thus captures the general methodology shared across this family of attacks.
GCG and similar methods craft adv by searching for token sequences following the affirmation objective—maximizing the likelihood of an affirmative response for a given instruction (affirm; e.g., “Sure, here’s how to build a bomb”). This builds on the observation that prefilling the response with an affirmative prefix (prefilling attacks; prefill-onllama3jailbreak2024) often induces successful jailbreaks qi2024safetyalignmentjusttokens. To increase universality across instructions on the targeted model, GCG can be further optimized against multiple harmful instructions zou2023gcg-universaltransferableadversarialattacks, a common though computationally intensive strategy thompson2024flrtfluentstudentteacherredteaming; sadasivan2024fastadversarialattackslanguage. To improve jailbreak fluency and facilitate optimization without access to model weights, BEAST sadasivan2024fastadversarialattackslanguage builds upon GCG by narrowing the search space to suffixes of high likelihood under the targeted language model.
@@ 3.1 Experimental Setup
Our main analysis uses Gemma2-2B-it gemmateam2024gemma2improvingopen to enable scale and depth, with critical evaluations validated on Qwen2.5-{0.5B,1.5B,32B}-Instruct qwen2025qwen25technicalreport and Llama-3.1-8B-Instruct grattafiori2024llama3herdmodels.
We use GCG (default hyperparameters) to craft 1200 adversarial suffixes on Gemma2-2B-it, each optimized against a single behavior sampled from AdvBench zou2023gcg-universaltransferableadversarialattacks, following GCG’s widely used affirmation objective, integrated in other suffix-based jailbreaks sadasivan2024fastadversarialattackslanguage; hayase2024querybasedadversarialpromptgeneration; andriushchenko2025PRSjailbreakingleadingsafetyalignedllms. Combined with 741 harmful instructions from AdvBench and StrongReject souly2024strongrejectjailbreaks, these yield nearly 900K GCG jailbreak prompts of varying success, used throughout the paper. We also generate 30 BEAST suffixes, and 100 GCG suffixes each for Qwen2.5 models and for Llama3.1, for additional evaluation. We defer more technical details to the appendix (App. A.1).
We measure jailbreak success using StrongReject’s fine-tuned classifier souly2024strongrejectjailbreaks, which assigns a grade , with higher values indicating more effective jailbreaks. We label attack samples as successful (), failed (), or borderline (otherwise), based on the classifier’s grading. A suffix’s universality score is defined as its success rate across the full set of harmful instructions (w.r.t. a single, inspected model). Throughout the paper, we randomly sample adversarial suffixes for evaluations while diversifying their universality level.
@@ 3.2 Characterizing GCG Suffixes
We analyze over 1K single-instruction GCG suffixes on Gemma2 (for Qwen2.5-1.5B and Llama3.1 see App. B.1) and reveal that (a) GCG suffixes show varying levels of efficacy, and (b) they often generalize beyond their explicit affirmation objective.
First, single-instruction GCG suffixes often generalize beyond their target instruction, exhibiting varying universality, a phenomenon also noted in contemporary work by huang-etal-2025-stronger. As Fig. 2(a) shows, most suffixes generalize to multiple tested harmful instructions, with the strongest succeeding in 20–60% of cases. A similar trend is observed with BEAST jailbreak suffixes (Fig. 11).
@@
These observations motivate our central question: what underlying mechanisms enable the effectiveness of different GCG suffixes, and particularly the emergent strong, universal suffixes?
@@ 4.1 Localizing the Critical Information Flow
Aiming to localize the critical information flow from the adversarial suffix (adv), we perform attention knockout gevaDissectingRecallFactual2023, as it is the sole component enabling information to transfer across token representations elhage2022solu.
Experimental setting. We sample 1K successful jailbreaks across suffixes of diverse universality, and perform attention knockout on each edge departing from adv to the following token subsequences (Fig. 1), by masking the edge’s attention (i.e., setting its attention logits to , in all layers). Then, we measure the Jailbreak Flip Rate (JFR): the fraction of attacks that are flipped by knockout from success to failure. Higher JFR indicates greater edge importance in generating a successful jailbreak.
Knockout adv. Fig. 3(a) shows the JFR of different suffixes, for each edge. We find advchat to be overwhelmingly critical for the jailbreak; knocking out advchat consistently fails the attack (causes refusal), whereas other edges (e.g., advaffirm) only occasionally do so. Notably, the removal of advchat prevents the jailbreak from manipulating the model into starting the generation with an affirmative token (e.g., “Sure”), which, as observed by qi2024safetyalignmentjusttokens, may by itself fail the jailbreak. To rule out this case, we perform a series of ablation studies on advchat’s knockout, forcing the generation to start with various dummy tokens (e.g., white spaces, additional sequence of chat tokens, and random punctuation), as well as an affirmative token (e.g., Sure), and find that the strong trend persists (JFR ; Fig. 13, App. B.2).
Knockout adv, under prefilling. Given the previous results (Fig. 3(a)), it is possible that the role of advchat is primarily to supply the affirmative response prefix; thus, failing the affirmation implies failing the jailbreak. In what follows, we rule this out. We repeat the knockout, this time applying it under prefilling, that is, starting the generation after an affirmative response prefix (§3.1). As Fig. 3(b) demonstrates, advchat still has the highest JFR among edges, with most suffixes having JFR (i.e., generally, this edge’s knockout mainly fails the prefilled attack). This shows the suffixes’ critical role extends beyond naïvely inducing affirmation.
@@ 4.2 Restoring Jailbreak via Shallow Patching
Experiment setting. To isolate the causal effect of the information flow to chat on the jailbreak behavior, we perform a causal mediated analysis vig2020causalmediationanalysisinterpreting, by patching chat’s attention output activations wang2022interpretabilitywildcircuitindirect; zhang2024bestpracticesactivationpatching.222Borrowing vig2020causalmediationanalysisinterpreting’s terms, we test the indirect effect of the input prompt (specifically, of adv) on the jailbreak behavior (i.e., a property of the output), while viewing chat as the inspected mediator. For 300 instruction-matched pairs of failed attacks (either random or GCG suffixes) and successful attacks, we take the failed sample and patch its attention outputs at chat (all layers) with those from the successful counterpart. Namely, we form each patched example by retaining the failed prompt (including adv) and injecting the successful sample’s chat activations; if this restores the jailbreak, we attribute it to the transferred advchat pathway. To control the patch depth, we incrementally extend the patch to i tokens after chat (denoted chat+i).
Patching chat+i. Fig. 4 shows that patching only chat restores the jailbreak behavior in the majority of cases. This effect is slightly enhanced when increasing the depth of the patched token subsequence into the generation (i.e., i in chat+i), with the majority of the effect taking place only a few tokens deep, suggesting the attacks’ key mechanism is shallow.